OpenStudy (anonymous):

For which operations is the set {–1, 0, 1} closed? Choose all answers that are correct. A. addition B. division C. multiplication D. subtraction

OpenStudy (anonymous):

whoever gets it ill give a medal

OpenStudy (anonymous):

C. multiplcation

OpenStudy (anonymous):

thanks can you help me with the next one

OpenStudy (anonymous):

7 ÷ 1/3 = ___ what dose this equal

OpenStudy (anonymous):

yeah sure

OpenStudy (anonymous):

Which equations show that the set of whole numbers is closed under addition? Choose all answers that are correct. A. 0 + (–1) = –1 B. 1 + 1 = 2 C. 2 + 0 = 2 D. 2 + (–1) = 1

OpenStudy (anonymous):

I would say D but im not 100% sure